OpenStudy (anonymous):

What are the first three terms of the sequence: a_1 = 3 and a_n = 2(_an - 1)^2? a) 2, 8, 18 b) 3, 18, 648 c) 3, 32, 50 d) 2, 6, 12, 24

OpenStudy (amistre64):

a{1}=3 ; a{n+1} = 2(a{n})^2 a{2} = 2(3)^2 = 18 .... its gonna be b